OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 54. Find y when x = 8 im not sure if its 384 or 144 y varies directly as the square of x. When x = 3, y = 54. Find y when x = 8 im not sure if its 384 or 144 @Mathematics

jimthompson5910 (jim_thompson5910):

"y varies directly as the square of x" means that y = kx^2 When x = 3, y = 54, this mean is that 54 = k(3)^2 54 = k(9) 54 = 9k 54/9 = k 6 = k k = 6 So the equation is y = 6x^2 Now when x = 8, y is... y = 6x^2 y = 6(8)^2 y = 6(64) y = 384 So when x = 8, y = 384